# Benford behavior resulting from stick and box fragmentation processes

### v2

@@ 3.2. PDF of $Y^{(N)}$
$\frac{d}{dy}\mathcal{E}_{Y^{(N)}}(y)$ $\ =\ \int_{-C\sqrt{N}}^{\frac{y}{d}}\prod_{j=2}^{d-1}\int_{z_{m+(j-1)-d}}^{\frac{y-\sum_{i=1}^{j-1}z_{m+i-d}}{d-j+1}}e_{Y^{(N)}}\left(z_{m+1-d},\dots,z_{m-1},y-\sum_{i=1}^{d-1}z_{m+i-d}\right)dz_{m-1}$
$\quad\quad\cdots dz_{m+1-d}.$ (3.19)
3.3. Outline of problem
We are ready to formulate our problem. We want to show that $\mathfrak{m}_{d}^{(N)}$ converges to strong Benford behavior, which is equivalent to showing that $\log_{B}(\mathfrak{m}_{d}^{(N)})$ converges to being equidistributed mod 1. Hence, in terms of the probability density function $f_{Y^{(N)}}(y)$ of $Y^{(N)}:=\log_{B}(\mathfrak{m}_{d}^{(N)})/\sqrt{N}$, we can formulate our problem as to showing
$F_{N}(a,b)\ :=\ \sum_{n=-dC\cdot N}^{dC\cdot N-1}\int_{\frac{a+n}{\sqrt{N}}}^{\frac{b+n}{\sqrt{N}}}f_{Y^{(N)}}(y)dy\ \approx\ b-a,$ (3.20)
for all $(a,b)\subset(0,1)$ for sufficiently large $N$. Let the main term and the error term of $F_{N}(a,b)$ be defined respectively as
$\mathcal{M}_{N}(a,b)$ $\ :=\ \sum_{n=-dC\cdot N}^{dC\cdot N-1}\int_{\frac{a+n}{\sqrt{N}}}^{\frac{b+n}{\sqrt{N}}}\frac{d}{dy}\mathcal{M}_{Y^{(N)}}(y)dy;$ (3.21)
$\mathcal{E}_{N}(a,b)$ $\ :=\ \sum_{n=-dC\cdot N}^{dC\cdot N-1}\int_{\frac{a+n}{\sqrt{N}}}^{\frac{b+n}{\sqrt{N}}}\frac{d}{dy}\mathcal{E}_{Y^{(N)}}(y)dy.$ (3.22)
@@ 3.3. Outline of problem
$\mathcal{M}_{N}(a,b)\ \approx\ b-a,\quad\mathcal{E}_{N}(a,b)\ \approx\ 0.$ (3.23)
3.4. Upper bound on error term $\mathcal{E}_{N}(a,b)$
We want to show that the error term $\mathcal{E}_{N}(a,b)$ is negligible. We start by giving some bounds on $d\mathcal{E}_{Y^{(N)}}(y)/dy$ from (3.2). Recall that $e_{Y^{(N)}}(z_{m+1-d},\dots,z_{m})$ is the error term of the joint density function (3.1). Let $\ell$] be short hand for $\{1,\dots,\ell\}$. Then
$\frac{1}{C_{m}^{m+1-d}}e_{Y^{(N)}}(z_{m+1-d},\dots,z_{m})$
@@ 3.4. Upper bound on error term $\mathcal{E}_{N}(a,b)$
$\quad\ =\ \left(\Phi(z_{m+1-d})^{m-d}+B(z_{m+1-d})\right)\left(\sum_{\begin{subarray}{c}S\subset[d]\\ S\neq d\end{subarray}}\left(\prod_{i\in S}\varphi(z_{m+i-d})\right)\left(\prod_{i\in[d]\setminus S}A(z_{m+i-d})\right)\right)$
@@
$\ \ll\ \sum_{\begin{subarray}{c}S\subset[d]\\ S\neq d\end{subarray}}\left(\prod_{i\in S}\varphi(z_{m+i-d})\right)\left(\prod_{i\in[d]\setminus S}A(z_{m+i-d})\right)+N^{-\delta}\prod_{i=1}^{d}\varphi(z_{m+i-d}).$ (3.24)
We know from (3.2) that to bound $d\mathcal{E}_{Y^{(N)}}(y)/dy$, it suffices to bound
$D_{1}(y)$ $\ :=\ \sum_{\begin{subarray}{c}S\subset[d]\\ S\neq[d]\end{subarray}}\int_{-C\sqrt{N}}^{\frac{y}{d}}\prod_{j=2}^{d-1}\int_{z_{m+(j-1)-d}}^{\frac{y-\sum_{i=1}^{j-1}z_{m+i-d}}{d-j+1}}\left(\prod_{i\in S}\varphi(z_{m+i-d})\right)\left(\prod_{i\in[d]\setminus S}A(z_{m+i-d})\right)\Bigg{|}_{z_{m}=y-\sum_{i=1}^{d-1}z_{m+i-d}}$
$\quad\quad dz_{m-1}\cdots dz_{m+1-d}$ (3.25)
@@
$D_{2}(y)\ :=\ N^{-\delta}\int_{-C\sqrt{N}}^{\frac{y}{d}}\prod_{j=2}^{d-1}\int_{z_{m+(j-1)-d}}^{\frac{y-\sum_{i=1}^{j-1}z_{m+i-d}}{d-j+1}}\prod_{i=1}^{d}\varphi(z_{m+i-d})\Bigg{|}_{z_{m}=y-\sum_{i=1}^{d-1}z_{m+i-d}}dz_{m-1}\cdots dz_{m+1-d},$ (3.26)
because then we have $d\mathcal{E}_{Y^{(N)}}dy\ll D_{1}(y)+D_{2}(y)$. We first look at $D_{1}(y)$. Suppose that $k_{S}$ is the largest index such that $k\in[d]\setminus S$. Such an index $k$ exists, since $S\neq[d]$. First, assume that we have $S$’s where $k_{S}=d$. Then the integrand of (3.4) becomes
@@
$\quad\ \ll\ N^{-\frac{1}{2}-\delta}\left(\prod_{i\in S}\varphi(z_{m+i-d})\right)\left(\prod_{i\in[d-1]\setminus S}A(z_{m+i-d})\right),$ (3.27)
where in the last line we use the fact that $A(x)=O(N^{-1/2-\delta})$. Note that the lower and upper bounds on the interval of integration for each integral in (3.4) are $O(\sqrt{N})$, since $y,z_{m+1-d},\dots,z_{m-1}$ are all $O(\sqrt{N})$. Choose a constant $C^{\prime}$ such that $C^{\prime}\sqrt{N}$ is an upper bound on the upper bounds on the interval of integration for all the integrals in (3.4) and $C^{\prime}\sqrt{N}$ is a lower bound on the interval of integration for all the integrals in (3.4). Hence,
$\int_{-C\sqrt{N}}^{\frac{y}{d}}\prod_{j=2}^{d-1}\int_{z_{m+(j-1)-d}}^{\frac{y-\sum_{i=1}^{j-1}z_{m+i-d}}{d-j+1}}\left(\prod_{i\in S}\varphi(z_{m+i-d})\right)\left(\prod_{i\in[d]\setminus S}A(z_{m+i-d})\right)\Bigg{|}_{z_{m}=y-\sum_{i=1}^{d-1}z_{m+i-d}}$
@@
$\quad\ \ll\ N^{-1/2-\delta}\underbrace{\int_{-C^{\prime}\sqrt{N}}^{C^{\prime}\sqrt{N}}\cdots\int_{-C^{\prime}\sqrt{N}}^{C\sqrt{N}}}_{d-1}\left(\prod_{i\in S}\varphi(z_{m+i-d})\right)\left(\prod_{i\in[d-1]\setminus S}N^{-1/2-\delta}\right)dz_{m-1}\cdots dz_{m+1-d}$
$\quad\ \ll\ N^{-1/2-\delta-|[d-1]\setminus S|\delta}\prod_{i\in S}\left(\int_{-C^{\prime}\sqrt{N}}^{C^{\prime}\sqrt{N}}\varphi(z_{m+i-d})dz_{m+i-d}\right)\ \ll\ N^{-1/2-(1+|[d-1]\setminus S|)\delta}\ \ll\ N^{-1/2-\delta},$ (3.28)
where in the last line we use the fact that
$\int_{-C^{\prime}\sqrt{N}}^{C^{\prime}\sqrt{N}}\varphi(x)dx\ \ll\ \int_{-\infty}^{\infty}\varphi(x)dx\ =\ 1,$ (3.29)
which we shall use frequently without reference from now on. Since the number of subsets $S$ of $[d]$ is finite, then (3.4) gives us
$D_{1}(y)\ \ll\ N^{-1/2-\delta}.$ (3.30)
The case when $k_{S}<d$ is similar, though more care is needed to identify the $O(N^{-1/2-\delta})$ decay, which essentially comes from the $A(z_{m+k_{S}-d})$ term. With this in mind, using the same method for the case when $k_{S}=d$ as well as the convolution formula for Gaussian PDF, we can again show that $D_{1}(y)\ll N^{-1/2-\delta}$. We leave the details to Appendix B.
We now turn to $D_{2}(y)$ as defined in (3.26), which is relatively more straightforward to bound. We have
$D_{2}(y)\ \ll\ N^{-\delta}\underbrace{\int_{-\infty}^{\infty}\cdots\int_{-\infty}^{\infty}}_{d-1}\prod_{i=1}^{d-1}\varphi(z_{m+i-d})\varphi\left(y-\sum_{i=1}^{d-1}z_{m+i-d}\right)dz_{m-1}\cdots dz_{m+1-d}.$ (3.31)

### v3

@@ 3.2 PDF of $Y^{(N)}$
$\quad\ =\ \int_{-\infty}^{\frac{y}{d}}\prod_{j=2}^{d-1}\int_{z_{m+(j-1)-d}}^{\frac{y-\sum_{i=1}^{j-1}z_{m+i-d}}{d-j+1}}f_{Z^{(N)}_{(m+1-d)},\dots,Z^{(N)}_{(m)},\textup{ error}}\left(z_{m+1-d},\dots,z_{m-1},y-\sum_{i=1}^{d-1}z_{m+i-d}\right)$
$\quad\quad\quad dz_{m-1}\cdots dz_{m+1-d}$
We would like to make some truncation to the tails of integrals in (3.2) and (3.2). The truncation method we now describe applies to both $dF_{Y^{(N)}}(y)/dy$ and $dF_{Y^{(N)},\textup{ main}}(y)/dy$, so we only demonstrate it for $dF_{Y^{(N)}}(y)/dy$. We want to show that
$\int_{-\infty}^{-N^{\delta^{\prime}}}\prod_{j=2}^{d-1}\int_{z_{m+(j-1)-d}}^{\frac{y-\sum_{i=1}^{j-1}z_{m+i-d}}{d-j+1}}f_{Z^{(N)}_{(m+1-d)},\dots,Z^{(N)}_{(m)}}\left(z_{m+1-d},\dots,z_{m-1},y-\sum_{i=1}^{d-1}z_{m+i-d}\right)$
$dz_{m-1}\cdots dz_{m+1-d}$ (89)
is asymptotically small for any $\delta^{\prime}$ satisfying $0<\delta^{\prime}<\delta$. By Proposition 5,
$f_{Z^{(N)}_{(m+1-d)},\dots,Z^{(N)}_{(m)}}\left(z_{m+1-d},\dots,z_{m-1},y-\sum_{i=1}^{d-1}z_{m+i-d}\right)$
$\quad\ =\ C_{m}^{m+1-d}f_{Z^{(N)}}(z_{m+1-d})\cdots f_{Z^{(N)}}(z_{m-1})f_{Z^{(N)}}\left(y-\sum_{i=1}^{d-1}z_{m+i-d}\right)(F_{Z^{(N)}}(z_{m+1-d}))^{m-d}.$ (90)
By Lemma 4, we know that $f_{Z^{(N)}}(z)=\varphi(z)+O(N^{-1/2-\delta})$, and $F_{Z^{(N)}}(z)=\Phi(z)+O(N^{-\delta})$, where $\varphi(z)=1/\sqrt{2\pi}e^{-z^{2}/2}$ is bounded between 0 and $1/\sqrt{2\pi}$ and $0\leq\Phi(z)\leq 1$, both of which are $O(1)$. Then,
$f_{Z^{(N)}_{(m+1-d)},\dots,Z^{(N)}_{(m)}}\left(z_{m+1-d},\dots,z_{m-1},y-\sum_{i=1}^{d-1}z_{m+i-d}\right)\ \ll\ f_{Z^{(N)}}(z_{m+1-d})\cdots f_{Z^{(N)}}(z_{m-1}).$ (91)
Returning to (3.2), we have that
$\int_{-\infty}^{-N^{\delta^{\prime}}}\prod_{j=2}^{d-1}\int_{z_{m+(j-1)-d}}^{\frac{y-\sum_{i=1}^{j-1}z_{m+i-d}}{d-j+1}}f_{Z^{(N)}_{(m+1-d)},\dots,Z^{(N)}_{(m)}}\left(z_{m+1-d},\dots,z_{m-1},y-\sum_{i=1}^{d-1}z_{m+i-d}\right)$
$dz_{m-1}\cdots dz_{m+1-d}$
$\quad\ \ll\ \int_{-\infty}^{-N^{\delta^{\prime}}}\underbrace{\int_{-\infty}^{\infty}\cdots\int_{-\infty}^{\infty}}_{d-2}f_{Z^{(N)}}(z_{m+1-d})f_{Z^{(N)}}(z_{m+2-d})\cdots f_{Z^{(N)}}(z_{m-1})$
$\quad\quad\quad dz_{m-1}\cdots dz_{m+2-d}dz_{m+1-d}$
$\quad\ =\ \int_{-\infty}^{-N^{\delta^{\prime}}}f_{Z^{(N)}}(z_{m+1-d})dz_{m+1-d}\prod_{i=2}^{d-1}\left(\int_{-\infty}^{\infty}f_{Z^{(N)}}(z_{m+i-d})dz_{m+i-d}\right).$ (92)
Since $f_{Z^{(N)}}(z)$ is a PDF, then $\int_{-\infty}^{\infty}f_{Z^{(N)}}(z)dz=1$. Moreover,
$\int_{-\infty}^{-N^{\delta^{\prime}}}f_{Z^{(N)}}(z_{m+1-d})dz_{m+1-d}$ $\ =\ \mathbb{P}\left(Z^{(N)}\leq-N^{\delta^{\prime}}\right)$
$\ \leq\ \mathbb{P}\left(|Z^{(N)}-\mathbb{E}[Z^{(N)}]|\geq N^{\delta^{\prime}}+\mathbb{E}[Z^{(N)}]\right)$
$\ \leq\ \frac{\textup{Var}[Z^{(N)}]}{(N^{\delta^{\prime}}+\mathbb{E}[Z^{(N)}])^{2}}\ =\ O(N^{-2\delta^{\prime}}),$ (93)
where the last line follows from Chebyshev’s inequality and the fact that the mean and the variance of $Z^{(N)}$ are 0 and $1$ respectively by construction (50) and (51). Hence, we have established that the tail of $dF_{Y^{(N)}}(y)/dy$ is $O(N^{-2\delta^{\prime}})$, i.e.,
$\frac{d}{dy}F_{Y^{(N)}}(y)$
$\quad\ =\ \int_{-N^{\delta^{\prime}}}^{\frac{y}{d}}\prod_{j=2}^{d-1}\int_{z_{m+(j-1)-d}}^{\frac{y-\sum_{i=1}^{j-1}z_{m+i-d}}{d-j+1}}f_{Z^{(N)}_{(m+1-d)},\dots,Z^{(N)}_{(m)}}\left(z_{m+1-d},\dots,z_{m-1},y-\sum_{i=1}^{d-1}z_{m+i-d}\right)$
$\quad\quad\quad dz_{m-1}\cdots dz_{m+1-d}+O(N^{-2\delta^{\prime}}),$ (94)
By the same truncation method, one can show that the tail of $dF_{Y^{(N)},\textup{ main}}(y)/dy$ is also $O(N^{-2\delta^{\prime}})$. The tail of $dF_{Y^{(N)},\textup{ error}}(y)/dy$ is the difference between the tail of $dF_{Y^{(N)}}(y)/dy$ and that of $dF_{Y^{(N)},\textup{ main}}(y)/dy$ by definition, then it is also $O(N^{-2\delta^{\prime}})$. Thus,
$\frac{d}{dy}F_{Y^{(N)},\textup{ main}}(y)$
$\quad\ =\ \int_{-N^{\delta^{\prime}}}^{\frac{y}{d}}\prod_{j=2}^{d-1}\int_{z_{m+(j-1)-d}}^{\frac{y-\sum_{i=1}^{j-1}z_{m+i-d}}{d-j+1}}f_{Z^{(N)}_{(m+1-d)},\dots,Z^{(N)}_{(m)},\textup{ main}}\left(z_{m+1-d},\dots,z_{m-1},y-\sum_{i=1}^{d-1}z_{m+i-d}\right)$
$\quad\quad\quad dz_{m-1}\cdots dz_{m+1-d}+O(N^{-2\delta^{\prime}}),$ (95)
$\frac{d}{dy}F_{Y^{(N)},\textup{ error}}(y)$
$\quad\ =\ \int_{-N^{\delta^{\prime}}}^{\frac{y}{d}}\prod_{j=2}^{d-1}\int_{z_{m+(j-1)-d}}^{\frac{y-\sum_{i=1}^{j-1}z_{m+i-d}}{d-j+1}}f_{Z^{(N)}_{(m+1-d)},\dots,Z^{(N)}_{(m)},\textup{ error}}\left(z_{m+1-d},\dots,z_{m-1},y-\sum_{i=1}^{d-1}z_{m+i-d}\right)$
$\quad\quad\quad dz_{m-1}\cdots dz_{m+1-d}+O(N^{-2\delta^{\prime}}).$ (96)
3.3 Outline of problem
We are ready to formulate Conjecture 1 in the language that we have set forth in Section 3.1. We want to show that $\mathfrak{m}_{d}^{(N)}$ converges to strong Benford behavior, which is equivalent to showing that $\log_{B}(\mathfrak{m}_{d}^{(N)})$ converges to being equidistributed mod 1. We know that $Y^{(N)}:=(\log_{B}(\mathfrak{m}_{d}^{(N)})-d\cdot N\mu_{P})/\sqrt{N}\sigma_{P}$. Since equidistribution mod 1 is invariant under translation [17], then it suffices to show that $\log_{B}(\mathfrak{m}_{d}^{(N)})-d\cdot N\mu_{P}$ converges to being equidistributed mod 1. Moreover, without loss of generality, we can assume $\sigma_{P}=1$ from now on, as the choice of renormalization would not change our final results. Hence, in terms of the PDF $f_{Y^{(N)}}(y)$, we can formulate our problem as to showing
$F_{N}(a,b)\ :=\ \mathbb{P}\left(Y^{(N)}\textup{ mod 1}\in(a,b)\right)\ =\ \sum_{n=-\infty}^{\infty}\int_{\frac{a+n}{\sqrt{N}}}^{\frac{b+n}{\sqrt{N}}}f_{Y^{(N)}}(y)dy\ \approx\ b-a,$ (97)
for all $(a,b)\subset(0,1)$ for sufficiently large $N$. Now, recall that we have truncated the outermost integral of $f_{Y^{(N)}}(y)=dF_{Y^{(N)}}(y)/dy$ at $N^{\delta^{\prime}}$ in Section 3.1, which is exactly the lower bound for $Z_{(m+1-d)}$. Since $Z_{(m+1-d)}\leq Z_{(m+2-d)}\leq\cdots\leq Z_{(m)}$, then the lower bound for their sum, $Y^{(N)}$, is $dN^{\delta^{\prime}}$. Hence, it is reasonable to truncate the sum of $F_{N}(a,b)$ at $n=\pm dN^{1/2+\delta^{\prime}}$ and instead show that
$\sum_{n=-dN^{1/2+\delta^{\prime}}}^{dN^{1/2+\delta^{\prime}}}\int_{\frac{a+n}{\sqrt{N}}}^{\frac{b+n}{\sqrt{N}}}f_{Y^{(N)}}(y)dy\ \approx\ b-a.$ (98)
We can make the truncation because the tails of the sum is negligible:
$\int_{\left|y\right|\geq dN^{\delta^{\prime}}}f_{Y^{(N)}}(y)dy$ $\ =\ \mathbb{P}\left(\left|Y^{(N)}\right|\geq dN^{\delta^{\prime}}\right)$
$\ \leq\ \mathbb{P}\left(\left|Y^{(N)}-\mathbb{E}[Y^{(N)}]\right|\geq dN^{\delta^{\prime}}-\mathbb{E}[Y^{(N)}]\right)$
$\quad\quad+\mathbb{P}\left(\left|Y^{(N)}-\mathbb{E}[Y^{(N)}]\right|\geq dN^{\delta^{\prime}}+\mathbb{E}[Y^{(N)}]\right)$
$\ \leq\ \frac{\textup{Var}[Y^{(N)}]}{(dN^{\delta^{\prime}}-\mathbb{E}[Y^{(N)}])^{2}}+\frac{\textup{Var}[Y^{(N)}]}{(dN^{\delta^{\prime}}+\mathbb{E}[Y^{(N)}])^{2}}\ =\ O(N^{-2\delta^{\prime}}),$ (99)
where the last line follows from Chebyshev’s inequality and the fact that the mean and the variance of $Y^{(N)}$ are finite, as it is the sum of $d$ random variables whose means and variances are finite. From now on, we shall only consider the bulk part of $F_{N}(a,b)$. Let the main term and the error term of $F_{N}(a,b)$ be defined respectively as
$\mathcal{M}_{N}(a,b)$ $\ :=\ \sum_{n=-dN^{1/2+\delta^{\prime}}}^{dN^{1/2+\delta^{\prime}}}\int_{\frac{a+n}{\sqrt{N}}}^{\frac{b+n}{\sqrt{N}}}\frac{d}{dy}F_{Y^{(N)},\textup{ main}}(y)dy;$ (100)
$\mathcal{E}_{N}(a,b)$ $\ :=\ \sum_{n=-dN^{1/2+\delta^{\prime}}}^{dN^{1/2+\delta^{\prime}}}\int_{\frac{a+n}{\sqrt{N}}}^{\frac{b+n}{\sqrt{N}}}\frac{d}{dy}F_{Y^{(N)},\textup{ error}}(y)dy.$ (101)
@@ 3.3 Outline of problem
$\mathcal{M}_{N}(a,b)\ \approx\ b-a,\quad\mathcal{E}_{N}(a,b)\ \approx\ 0.$ (102)
3.4 Upper bound on error term $\mathcal{E}_{N}(a,b)$
We first want to show that the error term $\mathcal{E}_{N}(a,b)$ is negligible. We start by giving some bounds on $dF_{Y^{(N)},\textup{error}}(y)/dy$ from (3.2). Recall that $f_{Z^{(N)}_{(m+1-d)},\dots,Z^{(N)}_{(m)},\textup{ error}}(z_{m+1-d},\dots,z_{m})$ is the error term of the joint density function (3.1). Let $\ell$] be short hand for $\{1,\dots,\ell\}$. Then
$\frac{1}{C_{m}^{m+1-d}}f_{Z^{(N)}_{(m+1-d)},\dots,Z^{(N)}_{(m)},\textup{ error}}(z_{m+1-d},\dots,z_{m})$
@@ 3.4 Upper bound on error term $\mathcal{E}_{N}(a,b)$
$\quad\ =\ \left(\Phi(z_{m+1-d})^{m-d}+B(z_{m+1-d})\right)\left(\sum_{S\subsetneq[d]}\left(\prod_{i\in S}\varphi(z_{m+i-d})\right)\left(\prod_{i\in[d]\setminus S}A(z_{m+i-d})\right)\right)$
@@
$\quad\ \ll\ \sum_{S\subsetneq[d]}\left(\prod_{i\in S}\varphi(z_{m+i-d})\right)\left(\prod_{i\in[d]\setminus S}A(z_{m+i-d})\right)+N^{-\delta}\prod_{i=1}^{d}\varphi(z_{m+i-d}).$ (103)
We know from (3.2) that to bound $dF_{Y^{(N)},\textup{ error}}(y)/dy$, it suffices to bound
$D_{1}(y)$ $\ :=\ \sum_{S\subsetneq[d]}\int_{-N^{\delta^{\prime}}}^{\frac{y}{d}}\prod_{j=2}^{d-1}\int_{z_{m+(j-1)-d}}^{\frac{y-\sum_{i=1}^{j-1}z_{m+i-d}}{d-j+1}}\left(\prod_{i\in S}\varphi(z_{m+i-d})\right)$
$\quad\quad\cdot\left(\prod_{i\in[d]\setminus S}A(z_{m+i-d})\right)\Bigg|_{z_{m}=y-\sum_{i=1}^{d-1}z_{m+i-d}}dz_{m-1}\cdots dz_{m+1-d}$ (104)
@@
$D_{2}(y)\ :=\ N^{-\delta}\int_{-N^{\delta^{\prime}}}^{\frac{y}{d}}\prod_{j=2}^{d-1}\int_{z_{m+(j-1)-d}}^{\frac{y-\sum_{i=1}^{j-1}z_{m+i-d}}{d-j+1}}\prod_{i=1}^{d}\varphi(z_{m+i-d})\Bigg|_{z_{m}=y-\sum_{i=1}^{d-1}z_{m+i-d}}dz_{m-1}\cdots dz_{m+1-d},$ (105)
because then by (3.2), we have
$\frac{d}{dy}F_{Y^{(N)},\textup{ error}}(y)\ \ll\ D_{1}(y)+D_{2}(y)+O(N^{-2\delta^{\prime}}).$ (106)
We first look at $D_{1}(y)$. Suppose that $k_{S}$ is the largest index such that $k\in[d]\setminus S$. Such an index $k$ exists, since $S\neq[d]$. First, consider $S$’s with $k_{S}=d$. Then the integrand of (3.4) becomes
@@
$\quad\ \ll\ N^{-\frac{1}{2}-\delta}\left(\prod_{i\in S}\varphi(z_{m+i-d})\right)\left(\prod_{i\in[d-1]\setminus S}A(z_{m+i-d})\right),$ (107)
where in the last line we use the fact that $A(x)=O(N^{-1/2-\delta})$. Now, recall the two truncations we performed in (3.2) and in (3.3). These truncations ensure that $z_{m+1-d}$ is at least $N^{\delta^{\prime}}$ and that $y$ is bounded above by $dN^{\delta^{\prime}}$ and below by $dN^{\delta^{\prime}}$. Since $y=z_{m+1-d}+\cdots+z_{m}$ and $z_{m+1-d}\leq\cdots\leq z_{m}$, then $z_{m}$ is at most $(2d-1)N^{\delta^{\prime}}$, which occurs when $y=dN^{\delta^{\prime}}$ and $z_{m+1-d}=\cdots=z_{m-1}=-N^{\delta^{\prime}}$. Hence, $y,z_{m+1-d},\dots,z_{m-1}$ are all $O(N^{\delta^{\prime}})$, and the lower and upper bounds on the domain of integration for each integral in (3.4) are $O(N^{\delta^{\prime}})$. Choose a constant $C$ such that $\pm CN^{\delta^{\prime}}$ are uniform upper and lower bounds on the domains of integration for all the integrals in (3.4). Hence,
$\int_{-N^{\delta^{\prime}}}^{\frac{y}{d}}\prod_{j=2}^{d-1}\int_{z_{m+(j-1)-d}}^{\frac{y-\sum_{i=1}^{j-1}z_{m+i-d}}{d-j+1}}\left(\prod_{i\in S}\varphi(z_{m+i-d})\right)\left(\prod_{i\in[d]\setminus S}A(z_{m+i-d})\right)\Bigg|_{z_{m}=y-\sum_{i=1}^{d-1}z_{m+i-d}}$
@@
$\quad\ \ll\ N^{-1/2-\delta}\underbrace{\int_{-CN^{\delta^{\prime}}}^{CN^{\delta^{\prime}}}\cdots\int_{-CN^{\delta^{\prime}}}^{CN^{\delta^{\prime}}}}_{d-1}\left(\prod_{i\in S}\varphi(z_{m+i-d})\right)\left(\prod_{i\in[d-1]\setminus S}N^{-1/2-\delta}\right)dz_{m-1}\cdots dz_{m+1-d}$
$\quad\ \ll\ N^{-1/2-\delta+|[d-1]\setminus S|(-1/2-\delta+\delta^{\prime})}\prod_{i\in S}\left(\int_{-CN^{\delta^{\prime}}}^{CN^{\delta^{\prime}}}\varphi(z_{m+i-d})dz_{m+i-d}\right)$
$\quad\ \ll\ N^{-1/2-\delta+|[d-1]\setminus S|(-1/2-\delta+\delta^{\prime})}\ \ll\ N^{-1/2-\delta},$ (108)
where in the last line we use the fact that $0<\delta^{\prime}<\delta$, as well as
$\int_{-CN^{\delta^{\prime}}}^{CN^{\delta^{\prime}}}\varphi(x)dx\ \ll\ \int_{-\infty}^{\infty}\varphi(x)dx\ =\ 1,$ (109)
which we shall use without reference for the remainder of the paper. Since there is only a finite number of subsets $S$ of $[d]$, then (3.4) has the estimate
$D_{1}(y)\ \ll\ N^{-1/2-\delta}.$ (110)
We move on to the case when $k_{S}<d$, which is similar to the $k_{S}=d$ case. Now, the $O(N^{-1/2-\delta})$ decay that we need comes from the $A(z_{m+k_{S}-d})$ term. With this in mind, using the same method for the case when $k_{S}=d$ as well as the convolution formula for Gaussian PDF, we can again show that $D_{1}(y)\ll N^{-1/2-\delta}$. We leave the details to B.
We now turn to $D_{2}(y)$ as defined in (105), which is relatively more straightforward to bound. We have
$D_{2}(y)\ \ll\ N^{-\delta}\underbrace{\int_{-\infty}^{\infty}\cdots\int_{-\infty}^{\infty}}_{d-1}\prod_{i=1}^{d-1}\varphi(z_{m+i-d})\varphi\left(y-\sum_{i=1}^{d-1}z_{m+i-d}\right)dz_{m-1}\cdots dz_{m+1-d}.$ (111)